# High-speed ditching of double curvature specimens with cavitation and ventilation

### v1

@@ 3.3 Effect of the horizontal speed
To gather more insight on the cavitation and ventilation modalities, especially in the final part of the impact phase, in Figure 15, the video frames at $\tau$=1 at all speeds are shown. It is possible to see that $U$=30 m/s at $\tau$=1 the cavitation region has expanded with respect to $\tau$=0.6, see Figure 13(c), but its rear edge has not moved significantly backwards. Therefore, this case can be ascribed to the incipient cavitation regime, also discussed in S3 (Iafrati & Grizzi, 2019). At $U$=35 m/s the cavitation region has reached the trailing edge and a ventilation region appears, which however remains confined around the trailing edge area, similarly to what happens in the intermediate regime, discussed above. Finally, at $U$=40 m/s and $U$=45 m/s $\tau$ =1 the ventilation front has clearly moved back towards the leading edge, therefore in these cases a cavitation-ventilation regime is observed. As a final comment, it can be observed that the threshold values that delineate the various regimes in S2 are analogous to those of S3.
As observed in (Iafrati, 2016), when varying the horizontal speed, the pressure should scale as $\rho\left(U^{2}+V^{2}\right)\,\sin{(\alpha+\beta)}$. As shown in 16, the scaling holds for the probes located at the front, whereas for the probes located at the rear the occurrence of cavitation introduces a lower bound for the pressures, which makes the scaling no longer valid. This is the reason why model scale tests could be reliable only if performed in de-pressurized conditions, which ensure the similarity of the cavitation number.
Figure 16: Time histories of the pressure scaled with the quantity $\rho\,(U^{2}+V^{2})$ for the tests of shape S2, at pitch angle 6 and different horizontal speeds.
@@
The normal force recorded by the front load cells $F_{zF}$ is more sensitive to the pressures acting on the front portion of the specimen, and for this reason it displays an increasing trend, which is consistent with the forward propagation of the spray root. The force intensity grows with speed, in line with the corresponding increase of the pressures. On the other hand, the rear load cells are more sensitive to the pressures acting on the back portion of the specimen and then, aside from the very early stage of the impact phase, they undergo the effects of the cavitation and ventilation phenomena. The load at the rear $F_{zR}$ exhibit a similar behaviour at the different speeds, assuming more and more negative values as the cavitation area expands up to about $\tau$= 0.5. Successively, when the ventilation starts, earlier for the test at $U$=45 m/s and then at 40 m/s, the force at the rear undergoes a sharp growth. It is worth noticing that this sharp variation of the loads introduces a significant displacement of the centre of loads during the water entry, which may have an important effect on the aircraft dynamics in a real ditching scenario. The total loads acting in the $z$- and $x$-directions are provided in the lower panels of Figure 17. The data indicate that the $x$-component is much smaller than the $z$-component and is characterized by a small hump occurring about $\tau$= 0.5. Although already visible in terms of the rear and forward components, the total $z$-force is characterized by a large increase with the horizontal speed, starting from the time at which the ventilation starts. Before that time, the positive force in the front part is partly balanced by the negative loads at the rear. When ventilation starts, there is a significant increase in the total loads due to the jump in the pressure values acting on the rear portion of the specimen. A positive contribution in this time interval also derives from the increase in the front force, as a consequence of the growth of the wetted area and to the forward propagation of the spray root, with the associated pressure peak.
By applying to the loads the same scaling proposed for the pressures in (Iafrati, 2016; Spinosa & Iafrati, 2021), the forces can be non-dimensionalized by the factor $\rho\,\left(U^{2}+V^{2}\right)\,S$ where $S$ is the area of the projection of the specimen on the $x-y$ plane. The non-dimensional forces plotted in Figure 18 indicate that the scaling holds for the forward component but not for the rear one, as a consequence of the occurrence of cavitation and ventilation.
Figure 18: Time histories of the forces scaled by the quantity $\rho\,(U^{2}+V^{2})\,S$ for the tests of shape S2, at pitch angle 6 and different horizontal speeds.
@@ 3.4 Effect of the pitch angle
Figure 21: Under-water frames comparison at $\tau$= 0.6 and 0.9 for the specimen S2 at $U$=40 m/s and at the pitch angles 4, 6 and 8.
@@
(d) $p$ line $l3$ Front probes - Pitch 4
(e) $p$ line $l3$ Front probes - Pitch 6
(f) $p$ line $l3$ Front probes - Pitch 8
@@
(j) $p$ line $l3$ Rear probes - Pitch 4
(k) $p$ line $l3$ Rear probes - Pitch 6
(l) $p$ line $l3$ Rear probes - Pitch 8
Figure 22: Front and rear pressure probe time histories for the specimen S2 at $U$=40 m/s and at the pitch angles 4, 6 and 8.
@@
As for the time histories recorded by the rear pressure probes, shown in Figure 22(g)-(l), it is shown that the drop to the vapour pressure values of all the probes located along the mid-line occurs over a shorter time period going from $\alpha$= 4 to $\alpha$= 6 and then $\alpha$= 8. Such a behaviour also is noticed at sensor line $l3$, where all pressure signals are delayed with respect to the midline due to the transverse curvature of the body. The pressure time histories also confirm that the duration of the cavitation phase is longer at 4, where ventilation starts just at the end of the impact phase, while at both 6 and 8 it starts at about $\tau=0.6$. It also appears that at 6 and 8, when ventilation occurs, the pressure rise to the atmospheric pressure is quite sharp. In contrast, at 4, the pressure growth is much milder. Such a pressure behaviour, in agreement with what discussed above regarding the video frames, is again similar to the intermediate regimes of cavitation, also observed for the shape S3 in Iafrati & Grizzi (2019).
@@ 3.5 Effects of the body curvatures
Figure 24: Under-water video frames of the specimens S1B, S2 and S3 during the water entry $U$=40 m/s and pitch angle 6 from $\tau$=0 to $\tau=0.8$. For the frames from $\tau$=0.2 to $\tau=0.6$ the edges of the geometric intersection area with the still water level are also plotted in blue. The direction of water entry is from bottom to the top.
@@
In this case, the front edge of the cavitation region appears to be situated in close proximity of the first contact point, whereas the rear edge does not exhibit a significant backward expansion in time. This situation is somewhat consistent with the incipient cavitation regime described in Section 3.3 and in Iafrati & Grizzi (2019). However, as it can be seen in Figure 25, at those times the leading edge of the specimen is already below the still water level, and a strong jet is formed at the front, which propagates towards the trailing edge. In this phase the trolley undergoes a sharp deceleration, and therefore the flow conditions are rather uncertain. For this reason the dynamics of the cavitation region for the specimen S1B is not discussed further.
@@
Figure 26: Front and rear pressure time histories for the three shapes, at $U$=40 m/s and at a pitch angle of 6.
@@
The difference between the the pressure fields of S2 and S3 on the bottom can also be visualized when comparing the iso-lines at $\tau$=0.3, 0.6 and 0.8 plotted for the shape S3 in Figure 27 with those of S2 shown above in Figure 9.
@@
(b) $\tau$= 0.6
(c) $\tau$= 0.8
Figure 27: Pressure iso-lines in the rear part of the specimen overlapped to the underwater video frames for the test of the shape S3 at U=40 m/s and at $\alpha=6^{\circ}$ at $\tau$=0.3, 0.6 and 0.8. The symmetry about the mid-line is exploited.
@@
Ventilation in S3 occurs in between $\tau$=0.5 and $\tau$=0.6. In Figure 28 the evolution of the process is visualised at six time instants within this non-dimensional time interval.
(a) Shape S2 - $\tau$=0.50
(b) Shape S2 - $\tau$=0.52
(c) Shape S2 - $\tau$=0.54
(d) Shape S2 - $\tau$=0.56
(e) Shape S2 - $\tau$=0.58
(f) Shape S2 - $\tau$=0.6
[…]
The mechanism of formation and development of the ventilation front in S3 is similar to S2, with a curved front starting from the centre at the trailing edge, which evolves very shortly into a straight one and reaches the point C.
@@
Figure 29: Time histories of the forces for the three shapes at $U$=40 m/s and pitch angle 6.
@@ 3.6 Cavitation zone propagation velocity
The estimated propagation velocities of the cavitation region $U_{\textrm{cav}}$, based on the image analysis and on the pressure time histories, are plotted in Figure 31 for the shapes S2 and S3 at 6, velocity ratio $V/U$=0.0375, and different horizontal speeds. It is worth noticing that in the graph supplementary test conditions with different horizontal velocities are considered, in addition to those listed in Table 2. $U_{\textrm{cav}}$ is plotted as a function of the horizontal velocity $U$, and linear fits of the data derived from the image analysis are also plotted, which both have a coefficient of determination $R^{2}$ very close to 1. It is worth noticing that the propagation velocity computed by using the pressure signals are very much in line with those derived from the image analysis.
@@ 3.7 Wetted area evolution
In the pioneering works of Von Karman (1929) and Wagner (1932) on vertical water entry, the growth of the wetted area is related to the time variation of the added mass and, therefore, to the impact loads. The primary distinction between the two theories is that, whereas von Karman’s theory considers solely the geometric intersection of the body with the free, the Wagner method also accounts for the water pile-up at the spray. In the work of Iafrati (2016) on the water entry of flat plates with a high horizontal speed, it was observed that the value of the pressure peak and the propagation of the spray root during the impact phase are related. Consequently, the evolution of the wetted area and the loads are also subject to a similar relationship. In light of the aforementioned discussion, it can be reasonably assumed that a comparable analogy can be expected in the front part of double-curvature specimens, where the pressures are found to be positive in all areas. Conversely, at the rear, the occurrence of negative pressures and, consequently, of cavitation and ventilation renders such an analogy invalid. In light of the aforementioned considerations, it is deemed insightful to conduct a more detailed investigation into the evolution of the wetted area in relation to the geometric intersection area.
@@
From the quantitative point of view it is possible to estimate the wetted area from the underwater video frames and to compare it with the geometrical intersection area. This is performed for the three shapes, using an image processing routine developed in MATLAB©. As shown in Figure 32, which refers to shape S3 at $\tau$=0.2 and 0.6, three Region Of Interests (ROIs) are manually selected and their areas are computed, first as pixel counts and then as the physical area in m2, by multiplying the pixel count for the pixel area derived from an image calibration. In particular, the defined ROIs are the front and rear wetted surface, which are delimited by the $x$-coordinate of the pressure probe row $r4$, the front and rear part of the geometric intersection area, still separated by the row $r4$, and finally the cavitation area. It is worth noticing that using this routine, the wetted areas are not exactly estimated rather their projection on the plane $X-Y$, parallel to the ground.
@@
[…]
@@ 4 Conclusions and Future Work
In addition to contributing to a deeper understanding of the physical phenomena occurring during the water entry at high horizontal speed and of the effects induced by changes in the body shape, the present paper provides a rather wide and detailed dataset, which could be used for an accurate validation of numerical approaches for the simulation of aircraft ditching or other water entry problems. Future work will focus on developing routine-based pattern recognition and machine learning algorithms to automate the selection and tracking in time of the wetted area, cavitation areas and the following ventilation front in underwater videos, which can be improved by also exploiting the information coming from the pressure probes.

### v2

@@ 3.3 Effect of the ditching speed
To gather more insight on the cavitation and ventilation modalities, especially in the final part of the impact phase, in Figure 13, the video frames at $\tau$=1 at all speeds are shown. It is possible to see that UU=30 m/s at $\tau$=1 the cavitation region has expanded with respect to $\tau$=0.6, see Figure 11(c), but its rear edge has not moved significantly backwards. Therefore, this case can be ascribed to the incipient cavitation regime, also discussed in S3 (Iafrati and Grizzi, 2019). At UU=35 m/s the cavitation region has reached the trailing edge and a ventilation region appears, which however remains confined around the trailing edge area, similarly to what happens in the intermediate regime, discussed above. Finally, at UU=40 m/s and UU=45 m/s $\tau$ =1 the ventilation front has clearly moved back towards the leading edge, therefore in these cases a cavitation-ventilation regime is observed. As a final comment, it can be observed that the threshold values that delineate the various regimes in S2 are analogous to those of S3.
Starting from this last consideration, this aspect can be related not only to the velocity values, but also to the cavitation number, which is defined as follows:
$Ca=\frac{p_{\textrm{atm}}-p_{\textrm{vap}}}{1/2\,\rho\,U^{2}}.$ (8)
The cavitation number, other relevant nondimensional parameters, and boundary layer characteristics for the tested shape conditions S2 and S3 are reported in Table 3 and Table 4, respectively. Table 4 includes additional test conditions that are not analysed in detail in this article but are discussed in Iafrati and Grizzi (2019).
$U$ [m/s] Ca [-] Cavitation Modality ReRe [-] $\delta$ [$m$] $C_{f}$ [-] $\ell_{v}$ [m]
29.87 0.222 Incipient $3.70\times 10^{7}$ 14.1 0.00242 $9.63\times 10^{-7}$
34.50 0.167 Intermediate $4.28\times 10^{7}$ 13.7 0.00237 $8.43\times 10^{-7}$
40.25 0.122 Cavit.+Vent. $4.99\times 10^{7}$ 13.2 0.00231 $7.31\times 10^{-7}$
46.19 0.093 Cavit.+Vent. $5.73\times 10^{7}$ 12.9 0.00226 $6.44\times 10^{-7}$
Table 3: Characteristic non-dimensional and boundary layer parameters for Shape S2
$U$ [m/s] Ca [-] Cavitation Modality ReRe [-] $\delta$ [$m$] $C_{f}$ [-] $\ell_{v}$ [m]
21.00 0.450 No Cavit. $2.60\times 10^{7}$ 15.1 0.00256 $1.33\times 10^{-6}$
26.80 0.276 No Cavit. $3.32\times 10^{7}$ 14.4 0.00246 $1.06\times 10^{-6}$
30.60 0.212 Incipient $3.79\times 10^{7}$ 14.0 0.00241 $9.42\times 10^{-7}$
34.50 0.167 Intermediate $4.28\times 10^{7}$ 13.7 0.00237 $8.43\times 10^{-7}$
35.70 0.156 Intermediate $4.43\times 10^{7}$ 13.6 0.00235 $8.17\times 10^{-7}$
37.20 0.143 Cavit.+Vent. $4.61\times 10^{7}$ 13.5 0.00234 $7.86\times 10^{-7}$
40.20 0.123 Cavit.+Vent. $4.98\times 10^{7}$ 13.2 0.00231 $7.32\times 10^{-7}$
45.20 0.097 Cavit.+Vent. $5.60\times 10^{7}$ 12.9 0.00227 $6.57\times 10^{-7}$
Table 4: CCharacteristic non-dimensional and boundary layer parameters for Shape S3
For both shapes no cavitation occurs for $\textrm{Ca}\geq$ 0.25. At Ca values around 0.20–0.22, incipient cavitation is observed, while for $\textrm{Ca}\leq$ 0.2 the cavitation region extends downstream, eventually leading to ventilation. This trend with Ca is consistent with the observations of Wang et al. (2001) and aligns with the description of the stages of sheet cavitation on hydrofoils discussed in Huang et al. (2019).
As a final note, it should be reminded that the effect of Reynolds number in water entry flows is generally considered secondary, as supported by Moghisi and Squire (1981); Facci et al. (2015). Additionally, the pressure drop at the rear can be explained solely by potential theory (Faltinsen, 2005; Newman, 1971; Semenov and Iafrati, 2006; Riccardi and Iafrati, 2004), indicating a negligible influence of viscous effects. Furthermore, data in Tables 3 and 4 show that both boundary layer thickness and viscous length, estimated approximately using the flat plate analogy do not vary significantly across the tested range of horizontal velocities. Therefore, boundary layer effects should remain essentially consistent across the cases and speeds under investigation.
@@
The normal force recorded by the front load cells $F_{zF}$ is more sensitive to the pressures acting on the front portion of the specimen, and for this reason it displays an increasing trend, which is consistent with the forward propagation of the spray root. The force intensity grows with speed, in line with the corresponding increase of the pressures. On the other hand, the rear load cells are more sensitive to the pressures acting on the back portion of the specimen and then, aside from the very early stage of the impact phase, they undergo the effects of the cavitation and ventilation phenomena. The load at the rear $F_{zR}$ exhibit a similar behaviour at the different speeds, assuming more and more negative values as the cavitation area expands up to about $\tau$= 0.5. Successively, when the ventilation starts, earlier for the test at UU=45 m/s and then at 40 m/s, the force at the rear undergoes a sharp growth. It is worth noticing that this sharp variation of the loads introduces a significant displacement of the centre of loads during the water entry, which may have an important effect on the aircraft dynamics in a real ditching scenario. The total loads acting in the $z$- and $x$-directions are provided in the lower panels of Figure 14. The data indicate that the $x$-component is much smaller than the $z$-component and is characterized by a small hump occurring about $\tau$= 0.5. Although already visible in terms of the rear and forward components, the total $z$-force is characterized by a large increase with the horizontal speed, starting from the time at which the ventilation starts. Before that time, the positive force in the front part is partly balanced by the negative loads at the rear. When ventilation starts, there is a significant increase in the total loads due to the jump in the pressure values acting on the rear portion of the specimen. A positive contribution in this time interval also derives from the increase in the front force, as a consequence of the growth of the wetted area and to the forward propagation of the spray root, with the associated pressure peak.
@@ 3.4 Effect of the pitch angle
Figure 17: Under-water frames comparison at $\tau$= 0.6 and 0.9 for the specimen S2 at UU=40 m/s and at the pitch angles 4, 6 and 8.
@@
(d) $p$ line $l3$ - Front probes - Pitch 4
(e) $p$ line $l3$ - Front probes - Pitch 6
(f) $p$ line $l3$ - Front probes - Pitch 8
@@
(j) $p$ line $l3$ - Rear probes - Pitch 4
(k) $p$ line $l3$ - Rear probes - Pitch 6
(l) $p$ line $l3$ - Rear probes - Pitch 8
Figure 18: Front and rear pressure probe time histories for the specimen S2 at UU=40 m/s and at the pitch angles 4, 6 and 8.
@@
As for the time histories recorded by the rear pressure probes, shown in Figure 18(g)-(l), it is shown that the drop to the vapour pressure values of all the probes located along the mid-line occurs over a shorter time period going from $\alpha$= 4° to $\alpha$= 6° and then $\alpha$= 8°. Such a behaviour also is noticed at sensor line $l3$, where all pressure signals are delayed with respect to the midline due to the transverse curvature of the body. The pressure time histories also confirm that the duration of the cavitation phase is longer at 4, where ventilation starts just at the end of the impact phase, while at both 6 and 8 it starts at about $\tau=0.6$. It also appears that at 6 and 8, when ventilation occurs, the pressure rise to the atmospheric pressure is quite sharp. In contrast, at 4, the pressure growth is much milder. Such a pressure behaviour, in agreement with what discussed above regarding the video frames, is again similar to the intermediate regimes of cavitation, also observed for the shape S3 in Iafrati and Grizzi (2019).
@@ 3.5 Effects of the body curvatures
Figure 20: Under-water video frames of the specimens S1B, S2 and S3 during the water entry UU=40 m/s and pitch angle 6 from $\tau$=0 to $\tau=0.8$. For the frames from $\tau$=0.2 to $\tau=0.6$ the edges of the geometric intersection area with the still water level are also plotted in blue. The direction of water entry is from bottom to the top.
@@
In this case, the front edge of the cavitation region appears to be situated in close proximity of the first contact point, whereas the rear edge does not exhibit a significant backward expansion in time. This situation is somewhat consistent with the incipient cavitation regime described in Section 3.3 and in Iafrati and Grizzi (2019). However, as it can be seen in Figure 21, at those times the leading edge of the specimen is already below the still water level, and a strong jet is formed at the front, which propagates towards the trailing edge. In this phase the trolley undergoes a sharp deceleration, and therefore the flow conditions are rather uncertain. For this reason the dynamics of the cavitation region for the specimen S1B is not discussed further.
@@
Figure 22: Front and rear pressure time histories for the three shapes, at UU=40 m/s and at a pitch angle of 6.
@@
The difference between the the pressure fields of S2 and S3 on the bottom can also be visualized when comparing pressure iso-lines which in Figure 23, are superimposed on the corresponding video frames at $\tau=0.3$ and $\tau=0.8$ and from $\tau=0.5$ to $\tau=0.6$ with a finer non-dimensional time resolution of $\Delta\tau=0.02$.
@@
(b) $\tau$=0.50
(c) $\tau$=0.52
(d) $\tau$=0.54
(e) $\tau$=0.56
(f) $\tau$=0.58
(g) $\tau$=0.6
(h) $\tau$= 0.8
Figure 23: Pressure iso-lines in the rear part of the specimens S3 overlapped to the underwater video frames for the water entry at U=40 m/s, at $\alpha=6^{\circ}$ at the specified $\tau$ values. The symmetry about the mid-line is exploited.
@@
Ventilation in S3 occurs in between $\tau$=0.5 and $\tau$=0.6. In Figure 23 the evolution of the process is visualised at six time instants within this non-dimensional time interval. The mechanism of formation and development of the ventilation front in S3 is similar to S2, with a curved front starting from the centre at the trailing edge, which evolves very shortly into a straight one and reaches the point C.
@@
Figure 24: Time histories of the forces for the three shapes at UU=40 m/s and pitch angle 6.
@@ 3.6 Cavitation zone propagation velocity
The estimated propagation velocities of the cavitation region $U_{\textrm{cav}}$, based on the image analysis and on the pressure time histories, are plotted in Figure 26 for the shapes S2 and S3 at 6, velocity ratio V/UV/U=0.0375, and different horizontal speeds. It is worth noticing that in the graph supplementary test conditions with different horizontal velocities are considered, in addition to those listed in Table 2. $U_{\textrm{cav}}$ is plotted as a function of the horizontal velocity $U$, and linear fits of the data derived from the image analysis are also plotted, which both have a coefficient of determination $R^{2}$ very close to 1. It is worth noticing that the propagation velocity computed by using the pressure signals are very much in line with those derived from the image analysis.
@@ 3.7 Wetted area evolution
In the pioneering works of Von Karman (1929) and Wagner (1932) on vertical water entry, the growth of the wetted area is related to the time variation of the added mass and, therefore, to the impact loads. The primary distinction between the two theories is that, whereas von Karman’s theory considers solely the geometric intersection of the body with the free, the Wagner method also accounts for the water pile-up at the spray. In the work of Iafrati (2016) on the water entry of flat plates with a high horizontal speed, it was observed that the value of the pressure peak and the propagation of the spray root during the impact phase are related. Consequently, the evolution of the wetted area and the loads are also subject to a similar relationship. In light of the aforementioned discussion, it can be reasonably assumed that a comparable analogy can be expected in the front part of double-curvature specimens, where the pressures are found to be positive in all areas. Conversely, at the rear, the occurrence of negative pressures and, consequently, of cavitation and ventilation renders such an analogy invalid. Given the aforementioned considerations, it is considered useful to conduct a more detailed investigation into the evolution of the wetted area in relation to the geometric intersection area.
@@
From the quantitative point of view it is possible to estimate the wetted area from the underwater video frames and to compare it with the geometrical intersection area. This is performed for the three shapes, using an image processing routine developed in MATLAB©. As shown in Figure 27, which refers to shape S3 at $\tau$=0.2 and 0.6, three Region Of Interests (ROIs) are manually selected and their areas are computed, first as pixel counts and then as the physical area in m2, by multiplying the pixel count for the pixel area derived from an image calibration. In particular, the defined ROIs are the front and rear wetted surface, which are delimited by the $x$-coordinate of the pressure probe row $r4$, the front and rear part of the geometric intersection area, still separated by the row $r4$, and finally the cavitation area. It is worth noticing that using this routine, the wetted areas are not exactly estimated rather their projection on the plane XYX-Y, parallel to the ground.
@@
It is observed that at the front both the wetted and the geometric intersection areas grow between $\tau$= 0.2 and $\tau$= 0.6, in particular the wetted area grows at a slightly higher rate. Owing to the much smoother curvature, specimen S1B is characterized by the lowest growth rates. As to S2 and S3, it can be noticed that in the interval $\tau\in(0.2\>0.6)$ the growth of the wetted area is about the same, despite the fact that S3 exhibits a higher growth rate of the geometric intersection area.
@@ 4 Conclusions and Future Work
In addition to contributing to a deeper understanding of the physical phenomena occurring during the water entry at high horizontal speed and of the effects induced by changes in the body shape, the present paper provides a rather wide and detailed dataset, which could be used for an accurate validation of numerical approaches for the simulation of aircraft ditching or other water entry problems. Future work will focus on developing routine-based pattern recognition algorithms to automate the selection and tracking in time of the wetted area, cavitation areas and the following ventilation front in underwater videos, which can be improved by also exploiting the information coming from the pressure probes.
Supplementary Material
Supplementary materials are available for download at: https://www.sciencedirect.com/science/article/pii/S0029801825018931#ecom0001
The graphical abstract can be viewed at this link: https://ars.els-cdn.com/content/image/1-s2.0-S0029801825018931-ga1_lrg.jpg